40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1108     Entry time: Mon Nov 3 19:12:27 2008
Author: alberto 
Type: Update 
Category: General 
Subject: Transverse mode spacing measurement for the X arm 
I know a lot of expectations have been building up on these days in the scientific community at the 40m towards a conclusive elog entry about the g-factor measurement of the X arm cavity.
The reason of the delay is that the results are still under review by the author. It turned out that the measurements of the transverse mode spacing have been performed on the beat
of the TEM02/20 and TEM00 modes between the two laser beams instead of on the beat between 00 and 01/10. However, the results posted on the elog in the last weeks seem likewise correct,
in particular my plot of the HOM of the sidebands.

Anyways, lately I have been trying to repeat the measurement on the beat of TEM01/10 with 00 but, despite all the efforts and the countless configurations tried (on the locking of
the arm, on the tilt of the mirrors, on the injection of the secondary beams, on the chopping with the blade), only the beat of TEM10 has been measured - although quite clearly -
whereas that of TEM01 has so far hidden itself.

The search continues but even if it does not succeeds, a summarizing document is going to be posted soon.

Here I attach a plot that shows the kind of difficulties trying to detect TEM10. The red neat peak is the beat of TEM01 whereas the other curves are some of the resulting
resonances after trying to couple TEM10 with 00 (or vice versa, according to whether I'm locking the cavity to the 00 mode of the main laser or to that of the secondary beam).
Attachment 1: 2008-11-02_summarizingplot.png  19 kB  | Hide | Hide all
2008-11-02_summarizingplot.png
ELOG V3.1.3-